# Reviving the Search for Indium-Based Superconductors: Theoretical Prediction of Semimetallic Superconductivity in Cubic Nd3In

### v2

@@ Abstract
The quest for materials that simultaneously exhibit superconductivity and non trivial topology has drawn significant attention in recent years, driven by their potential to host exotic quantum states. Their unique coexistence often leads to rich physics and potential applications in quantum technologies. Here, we predict cubic Nd3In as an exceptional candidate in this class, combining strong-coupling superconductivity with distinctive topological features. Using first-principles calculations, we find that the strong-coupling superconductivity in Nd3In arises primarily due to pronounced Fermi surface nesting, leading to an electron-phonon coupling constant of $\lambda=1.39$. Our fully anisotropic Migdal–Eliashberg analysis yields a superconducting transition temperature $T_{c}\approx 14$ K at ambient pressure, increasing to 18 K under 15 GPa. Beyond superconductivity, Nd3In is found to be a Weyl semimetal, as evidenced by the presence of Fermi arcs and nontrivial $\mathbb{Z}_{2}$ topological invariants, confirming its topological nature. The combination of strong-coupling superconductivity and nontrivial topological states makes Nd3In a promising candidate for quantum transport and topological quantum computation.
@@ II Methodology and Crystal structure
$T_{c}=\frac{\omega_{\log}}{1.2}\exp\left[-\frac{1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right],$ (1)
where $\lambda$ is the EPC constant, $\omega_{\log}$ is the logarithmic average of phonon frequencies, and $\mu^{*}$ is the Coulomb pseudopotential. The total EPC parameter is given by
@@
while the Eliashberg spectral function itself is defined as
@@
where $g^{\nu}_{mn}(\mathbf{k},\mathbf{q})$ are the electron-phonon matrix elements, and $\omega_{\nu\mathbf{q}}$ are the phonon frequencies. To further investigate the electronic structure, the Fermi surface nesting function $\zeta(\mathbf{q})$ is computed as
@@ III.2 Phonon dispersion and electron-phonon interaction
The phonon dispersion and projected phonon density of states (PHDOS) for cubic Nd3In are shown in Figure 6. The absence of imaginary frequencies throughout the Brillouin zone confirms the dynamical stability of the structure. The phonon spectrum of Nd3In can be divided into three distinct frequency regions. This distribution reflects the expected behavior that heavier atoms like Nd vibrate at lower frequencies than lighter atoms such as In. A closer inspection of the phonon dispersion reveals the presence of soft phonon modes, or Kohn anomalies, along specific high-symmetry directions particularly along the $\Gamma$–X–M and R–X paths. Generally, these soft phonon modes are indicative of strong electron-phonon coupling [91, 92, 93, 94]. All three acoustic branches as shown in the left panel of Figure 6, exhibit noticeable softening along these directions, with the sixth vibrational mode showing pronounced softening along R–X1. We have also examined the pressure dependence of this phonon softening, as shown in Figure A1 (in Appendix A). The phonon dispersion weighted by mode-resolved electron-phonon coupling strength and the corresponding spectral functions were computed at 10 GPa and 15 GPa pressure. The softening intensifies under applied pressure, suggesting a possible enhancement in the superconducting critical temperature.
@@
To investigate the origin of the Kohn anomalies at ambient pressure, we computed the Fermi surface nesting function $\zeta(\mathbf{Q})$ using the EPW code on a dense $k$-grid of $64\times 64\times 64$ and a $q$-grid of $32\times 32\times 32$, as shown in Figure 7. The nesting function exhibits pronounced peaks along high-symmetry directions where phonon softening is observed, indicating strong Fermi surface nesting. At the $\Gamma$ point, the prominent peak corresponds to the entire Fermi surface nesting onto itself, which lacks physical significance. Figure 7 shows that $\zeta(\mathbf{Q})$ exhibits notably high values along the $\Gamma\rightarrow$ X, $\Gamma\rightarrow$ X1 directions and as well as between the $\Gamma\rightarrow$ M path. As previously discussed, this indicates that parallel segments of the Fermi surface are connected by specific wave vectors $\mathbf{Q}$, in accordance with the features observed in the calculated Fermi surfaces shown in Figure 3(b). Such nesting enhances the electron-phonon interaction, leading to strong-coupling and giving rise to Kohn anomalies [95, 84, 96, 97, 98]. We then performed a quantitative evaluation of the electron–phonon coupling (EPC) strength. Initially, density-functional perturbation theory (DFPT) calculations on a coarse $4\times 4\times 4$ $\mathbf{q}$-mesh yielded a preliminary EPC constant of $\lambda=0.982$. To obtain a more accurate value, we employed the EPW code, which utilizes maximally localized Wannier functions (MLWFs) to interpolate the electron–phonon matrix elements over dense Brillouin zone grids, specifically a $48\times 48\times 48$ $\mathbf{k}$-mesh and a $24\times 24\times 24$ $\mathbf{q}$-mesh. To construct the MLWFs, we used initial projections on the Nd $d$ and In $p$ orbitals. The accuracy of the Wannierization was verified by comparing the Wannier interpolated band structure with the original DFT results, as shown in Figure A2 (in Appendix A). From the EPW calculations, we obtained a refined electron–phonon coupling (EPC) constant of $\lambda=1.394$, indicating strong coupling and suggesting that Nd3In is a promising candidate for phonon-mediated superconductivity. The dominant contributions to the total EPC arise from soft phonon modes, as illustrated in the left panel of Figure 6, where the mode-resolved coupling strengths $\lambda_{q\nu}$ are projected onto the phonon dispersion. A pronounced enhancement of $\lambda_{q\nu}$ is observed along the $\Gamma$–X–M and R–X1 directions, which coincide with regions of phonon softening. The right panel of Figure 6 shows the Eliashberg spectral function $\alpha^{2}F(\omega)$ and the cumulative EPC function $\lambda(\omega)$. Analysis of $\lambda(\omega)$ reveals that phonons below 65 cm-1 contribute approximately 62.1% of the total EPC, while intermediate-frequency modes between 65 and 90 cm-1 account for 20.8%. The remaining contribution comes from higher-frequency modes.
@@ III.3 Superconductivity
The temperature evolution of the superconducting energy gap $\Delta_{nk}$ is presented in Figure 8 (a). At 2 K, the gap ranges from 1.75 to 2.30 meV, reflecting its anisotropic nature. As temperature increases, the gap gradually closes and disappears around 14 K, which we identify as the true superconducting transition temperature. This value shows a 32.16% enhancement over the isotropic Allen–Dynes estimate, highlighting the significance of momentum-dependent electron-phonon coupling considered in the fully anisotropic Migdal–Eliashberg formalism, as previously mentioned. Notably, Nd3In exhibits the highest superconducting transition temperature among the AuCu3-type semimetallic superconductors. Moreover, $T_{c}$ further increases under pressure, reaching 16.5 K at 10 GPa and 18 K at 15 GPa, as shown in Figure A3 and A4 (in Appendix A). Thus, Nd3In can be classified as a low-$T_{c}$ superconductor and, as previously discussed, holds strong potential for applications in technologies such as MRI, SRF cavities, fusion reactors, and particle accelerators. Furthermore, the ratio $3.932\Delta_{0}/k_{B}T_{c}=$ 3.93 exceeds the BCS weak-coupling limit of 3.52, underscoring the strong-coupling character of superconductivity in Nd3In, which primarily originates from the pronounced Fermi surface nesting. These findings clearly demonstrate that the isotropic McMillan–Allen–Dynes approach underestimates $T_{c}$ in materials with pronounced EPC anisotropy, while the anisotropic Migdal–Eliashberg theory provides a more robust and detailed understanding of the superconducting mechanism.
@@ III.4 Semimetallic topology
Before concluding, we briefly discuss the compositional variant In3Nd, with the corresponding results summarized in the Appendix A for reference. In3Nd crystallizes in the AuCu3-type structure with a lattice constant of 4.721 Å without SOC (4.684 Å with SOC). We find that this material also exhibits semimetallic superconductivity. Orbital-projected band structures and projected density of states without and with SOC are shown in Figures A5 and A6, respectively. Figure A5 reveals several type-I Weyl points that are gapped out upon inclusion of SOC, resulting in a continuous gap between the highest valence band and the lowest conduction band. Notably, a Weyl point is observed precisely at the Fermi level along the $M$$\Gamma$ path. The semimetallic nature of In3Nd is further supported by the presence of Fermi arcs, as illustrated in Figure A7. From the superconducting perspective, In3Nd exhibits a low transition temperature of $T_{c}=0.72$ K, with a superconducting gap of 0.11 meV and an electron-phonon coupling constant $\lambda=0.435$. This significant decrease in $T_{c}$ relative to Nd3In is primarily attributed to a substantial decrease in the total electronic density of states at the Fermi level, which decreases from 6.19 to 1.51 states/eV due to the Nd–In site interchange. The mode-resolved electron-phonon coupling strength, Eliashberg spectral function, and electronic density of states for In3Nd are detailed in Figure A8. All figures related to In3Nd are provided in the Appendix for reference. Interestingly, this valency-exchanged phenomenon is also observed in the La–In system, where both La3In and In3La crystallize in the AuCu3-type structure. While La3In is a conventional superconductor [115], its compositional counterpart In3La exhibits semimetallic superconductivity [116, 117].
@@ IV Conclusions
Appendix
Figure A1: Phonon dispersion relations of Nd3In under different hydrostatic pressures. The fat bands in the dispersion plots indicate modes with strong electron–phonon coupling. (a) At ambient pressure ($P=0$ GPa), the phonon spectrum exhibits distinct acoustic and optical branches with relatively low-frequency modes. (b) At moderate pressure ($P=10$ GPa), noticeable softening of the acoustic phonon branches occurs, reflecting pressure-induced modifications in lattice dynamics. (c) At higher pressure ($P=15$ GPa), further softening is observed, particularly in the low-frequency region. Correspondingly, the electron-phonon coupling constant $\lambda$ increases with pressure, suggesting enhanced superconducting interactions.
Figure A2: Electronic band structure of Nd3In without spin–orbit coupling (SOC). The Wannier90-interpolated bands are shown as blue dashed lines, while the DFT reference band structure is displayed as solid black lines.
[…]
Figure A4: Pressure dependence of the superconducting transition temperature $T_{c}$ for Nd3In. The values of $T_{c}$ are extracted from fully anisotropic Eliashberg theory calculations at three different pressures: 0 GPa, 10 GPa, and 15 GPa. A clear enhancement of $T_{c}$ with increasing pressure is observed, highlighting the role of pressure in tuning electron-phonon interactions and enhancing superconducting behavior in Nd3In.
Figure A5: Orbital-projected band structure and corresponding density of states (PDOS) for In3Nd in the absence of spin–orbit coupling. Scatter point sizes correspond to the strength of each orbital’s contribution. At the Fermi level, the total DOS is low (1.51 states/eV), suggesting fewer available electronic states.
Figure A6: Band structure of In3Nd with orbital projections and projected density of states (PDOS). The size of the scatter points is proportional to the weight of each orbital’s contribution. The density of states at the Fermi level slightly decreases to approximately 1.38 states/eV upon inclusion of spin–orbit coupling (SOC). It can also be seen that several band gaps open due to SOC, as indicated in the figure.
Figure A7: Fermi arcs observed in the compositional variant In3Nd, providing evidence for its semimetallic character.
Figure A8: Phononic and superconducting properties of In3333Nd. The left panel shows the phonon dispersion weighted by the mode-resolved electron–phonon coupling strength $\lambda{\mathbf{q}\nu}$, where darker regions indicate stronger coupling. The middle panel presents the atom-projected phonon density of states (PDOS), and the right panel displays the Eliashberg spectral function $\alpha^{2}F(\omega)$ together with the integrated EPC function $\lambda(\omega)$, highlighting the superconducting characteristics.

### v3

# Reviving the Search for Indium-Based Superconductors: First-Principles Prediction of Semimetallic Superconductivity in Cubic Nd3In

@@ Abstract
The quest for materials that simultaneously exhibit superconductivity and non trivial topology has drawn significant attention in recent years, driven by their potential to host exotic quantum states. Their unique coexistence often leads to rich physics and potential applications in quantum technologies. Here, we predict cubic Nd3In as an exceptional candidate in this class, combining strong-coupling superconductivity with distinctive topological features. Using first-principles calculations, we find that the strong-coupling superconductivity in Nd3In arises primarily due to pronounced Fermi surface nesting, leading to an electron-phonon coupling constant of $\lambda=1.39$. Our fully anisotropic Migdal–Eliashberg analysis predicts a superconducting transition temperature $T_{c}\approx 14\ K$ at ambient pressure, which is the highest value reported so far among cubic semimetallic superconductors. When subjected to a pressure of 15 GPa, $T_{c}$ increases further to 18 K. Beyond superconductivity, Nd3In is found to be a Weyl semimetal, as evidenced by the presence of Fermi arcs and nontrivial $\mathbb{Z}_{2}$ topological invariants, confirming its topological nature. The combination of strong-coupling superconductivity and nontrivial topological states makes Nd3In a promising candidate for quantum transport and topological quantum computation.
@@ II Methodology and Crystal structure
$T_{c}=f_{1}f_{2}\frac{\omega_{\log}}{1.2}\exp\left[\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right],$ (1)
where $\omega_{\log}$ denotes the logarithmic average of phonon frequencies, and $\mu^{*}$ is the retarded Coulomb pseudopotential. The correction factors $f_{1}$ and $f_{2}$ are expressed as
$f_{1}f_{2}$ $=\left[1+\left(\frac{\lambda}{2.46(1+3.8\mu^{*})}\right)^{\frac{3}{2}}\right]^{\frac{1}{3}}$
×$\quad\times\left(1-\frac{\lambda^{2}(1-\omega_{2}/\omega_{\log})}{\lambda^{2}+3.312(1+6.3\mu^{*})^{2}}\right),$ (2)
with $\omega_{2}$ being the mean-square phonon frequency defined by
@@
The total electron-phonon coupling (EPC) constant $\lambda$ is given by
[…]
where $\alpha^{2}F(\omega)$ is the Eliashberg spectral function, defined as
@@
Here, $g^{\nu}_{mn}(\mathbf{k},\mathbf{q})$ are the electron-phonon matrix elements, and $\omega_{\nu\mathbf{q}}$ are the phonon frequencies. To further investigate the electronic structure, the Fermi surface nesting function $\zeta(\mathbf{q})$ is computed as
@@ III.2 Phonon dispersion and electron-phonon interaction
The phonon dispersion and projected phonon density of states (PHDOS) for cubic Nd3In are shown in Figure 6. The absence of imaginary frequencies throughout the Brillouin zone confirms the dynamical stability of the structure. The phonon spectrum of Nd3In can be divided into three distinct frequency regions. This distribution reflects the expected behavior that heavier atoms like Nd vibrate at lower frequencies than lighter atoms such as In. A closer inspection of the phonon dispersion reveals the presence of soft phonon modes, or Kohn anomalies, along specific high-symmetry directions particularly along the $\Gamma$–X–M and R–X paths. Generally, these soft phonon modes are indicative of strong electron-phonon coupling [91, 92, 93, 94]. All three acoustic branches as shown in the left panel of Figure 6, exhibit noticeable softening along these directions, with the sixth vibrational mode showing pronounced softening along R–X1. We have also examined the pressure dependence of this phonon softening, as shown in Figure S1 of the Supplementary Materials [95]). The phonon dispersion weighted by mode-resolved electron-phonon coupling strength and the corresponding spectral functions were computed at 10 GPa and 15 GPa pressure. The softening intensifies under applied pressure, suggesting a possible enhancement in the superconducting critical temperature.
@@
To investigate the origin of the Kohn anomalies at ambient pressure, we computed the Fermi surface nesting function $\zeta(\mathbf{Q})$ using the EPW code on a dense $k$-grid of $64\times 64\times 64$ and a $q$-grid of $32\times 32\times 32$, as shown in Figure 7. The nesting function exhibits pronounced peaks along high-symmetry directions where phonon softening is observed, indicating strong Fermi surface nesting. At the $\Gamma$ point, the prominent peak corresponds to the entire Fermi surface nesting onto itself, which lacks physical significance. Figure 7 shows that $\zeta(\mathbf{Q})$ exhibits notably high values along the $\Gamma\rightarrow$ X, $\Gamma\rightarrow$ X1 directions and as well as between the $\Gamma\rightarrow$ M path. As previously discussed, this indicates that parallel segments of the Fermi surface are connected by specific wave vectors $\mathbf{Q}$, in accordance with the features observed in the calculated Fermi surfaces shown in Figure 3(b). Such nesting enhances the electron-phonon interaction, leading to strong-coupling and giving rise to Kohn anomalies [96, 84, 97, 98, 99]. We then performed a quantitative evaluation of the electron–phonon coupling (EPC) strength. Initially, density-functional perturbation theory (DFPT) calculations on a coarse $4\times 4\times 4$ $\mathbf{q}$-mesh yielded a preliminary EPC constant of $\lambda=0.982$. To obtain a more accurate value, we employed the EPW code, which utilizes maximally localized Wannier functions (MLWFs) to interpolate the electron–phonon matrix elements over dense Brillouin zone grids, specifically a $48\times 48\times 48$ $\mathbf{k}$-mesh and a $24\times 24\times 24$ $\mathbf{q}$-mesh. To construct the MLWFs, we used initial projections on the Nd $d$ and In $p$ orbitals. The accuracy of the Wannierization was verified by comparing the Wannier interpolated band structure with the original DFT results, as shown in Figure S2 (see Supplementary Materials [95]). From the EPW calculations, we obtained a refined electron–phonon coupling (EPC) constant of $\lambda=1.394$, indicating strong coupling and suggesting that Nd3In is a promising candidate for phonon-mediated superconductivity. The dominant contributions to the total EPC arise from soft phonon modes, as illustrated in the left panel of Figure 6, where the mode-resolved coupling strengths $\lambda_{q\nu}$ are projected onto the phonon dispersion. A pronounced enhancement of $\lambda_{q\nu}$ is observed along the $\Gamma$–X–M and R–X1 directions, which coincide with regions of phonon softening. The right panel of Figure 6 shows the Eliashberg spectral function $\alpha^{2}F(\omega)$ and the cumulative EPC function $\lambda(\omega)$. Analysis of $\lambda(\omega)$ reveals that phonons below 65 cm-1 contribute approximately 62.1% of the total EPC, while intermediate-frequency modes between 65 and 90 cm-1 account for 20.8%. The remaining contribution comes from higher-frequency modes.
@@ III.3 Superconductivity
The temperature evolution of the superconducting energy gap $\Delta_{nk}$ is presented in Figure 8 (a). At 2 K, the gap ranges from 1.75 to 2.30 meV, reflecting its anisotropic nature. As temperature increases, the gap gradually closes and disappears around 14 K, which we identify as the true superconducting transition temperature. This value shows a 32.16% enhancement over the isotropic Allen–Dynes estimate, highlighting the significance of momentum-dependent electron-phonon coupling considered in the fully anisotropic Migdal–Eliashberg formalism, as previously mentioned. Notably, Nd3In exhibits the highest superconducting transition temperature among the AuCu3-type semimetallic superconductors. Moreover, $T_{c}$ further increases under pressure, reaching 16.5 K at 10 GPa and 18 K at 15 GPa, as shown in Figure S3 and S4 (in Supplementary Materials [95]). Thus, Nd3In can be classified as a low-$T_{c}$ superconductor and, as previously discussed, holds strong potential for applications in technologies such as MRI, SRF cavities, fusion reactors, and particle accelerators. Furthermore, the ratio $3.932\Delta_{0}/k_{B}T_{c}=$ 3.93 exceeds the BCS weak-coupling limit of 3.52, underscoring the strong-coupling character of superconductivity in Nd3In, which primarily originates from the pronounced Fermi surface nesting. These findings clearly demonstrate that the isotropic McMillan–Allen–Dynes approach underestimates $T_{c}$ in materials with pronounced EPC anisotropy, while the anisotropic Migdal–Eliashberg theory provides a more robust and detailed understanding of the superconducting mechanism.
@@ III.4 Semimetallic topology
Before concluding, we briefly discuss the compositional variant In3Nd, with the corresponding results summarized in Supplementary Materials [95]. In3Nd crystallizes in the AuCu3-type structure with a lattice constant of 4.721 Å without SOC (4.684 Å with SOC). We find that this material also exhibits semimetallic superconductivity. Orbital-projected band structures and projected density of states without and with SOC are shown in Figures S5 and S6, respectively. Figure S5 reveals several type-I Weyl points that are gapped out upon inclusion of SOC, resulting in a continuous gap between the highest valence band and the lowest conduction band. Notably, a Weyl point is observed precisely at the Fermi level along the $M$$\Gamma$ path. The semimetallic nature of In3Nd is further supported by the presence of Fermi arcs, as illustrated in Figure S7. From the superconducting perspective, In3Nd exhibits a low transition temperature of $T_{c}=0.72$ K, with a superconducting gap of 0.11 meV and an electron-phonon coupling constant $\lambda=0.435$. This significant decrease in $T_{c}$ relative to Nd3In is primarily attributed to a substantial decrease in the total electronic density of states at the Fermi level, which decreases from 6.19 to 1.51 states/eV due to the Nd–In site interchange. The mode-resolved electron-phonon coupling strength, Eliashberg spectral function, and electronic density of states for In3Nd are detailed in Figure S8. Interestingly, this valency-exchanged phenomenon is also observed in the La–In system, where both La3In and In3La crystallize in the AuCu3-type structure. While La3In is a conventional superconductor [116], its compositional counterpart In3La exhibits semimetallic superconductivity [117, 118].